# Ara-HOPE: Human-Centric Post-Editing Evaluation for Dialectal Arabic to Modern Standard Arabic Translation

### v1

@@ 4.2 Questionnaire Design
A structured Excel sheet was created to support annotation and data analysis. Each sheet contained three sets of columns, one per translation engine. Figure 1 shows two annotated examples for the Jais system. The first three columns show the DA sentence, the MSA gold translation, and the proposed machine translation. Columns 4-8 represent the five Ara-HOPE error categories, with annotators assigning severity scores from 0 (no error) to 2 (major error). Empty cells indicate a score of 0, and the final column sums the error scores for each sentence.
Figure 1: Example layout of the annotation questionnaire used for evaluating DA-MSA translations across fluency, meaning transfer, and adaptation error categories.
@@
Three physical copies of the Excel sheet were printed (one for each system) as annotators preferred working on paper. Each contained the same DA-MSA pairs, differing only in the translation engine output. Annotators marked scores manually, and data were later transferred back to the digital Excel template.
@@ 5.1 Inter Annotator Agreement
Inter-Annotator Agreement (IAA) measures consistency between evaluators, with high scores indicating a reliable and reproducible annotation framework. In our study, we employed Quadratic Weighted Cohen’s Kappa (QWK) because it effectively handles ordinal severity ratings by assigning greater penalties to larger disagreements. This makes it particularly suitable for error severity judgment on our 0-2 scale, where the magnitude of disagreement matters.
The IAA scores, as shown in Table 1 varied meaningfully across error categories: Fluency showed strong agreement for most MT systems, likely due annotators’ Arabic language expertise, which minimized inaccurate judgments. Meaning Transfer achieved robust agreement, likely due to the clear criteria regarding preservation of original meaning, as established in the Ara-HOPE taxonomy decision tree (Figure 5). Those definitions were particularly refined before the full evaluation during our pilot testing and feedback sessions. This consistency indicates a shared understanding of meaning preservation requirements among our annotators. Adaptation yielded the lowest agreement (0.122-0.28) due to its inherently subjective nature. Assessing contextual appropriateness of DA translations involves greater personal judgment than other error categories.
@@
Overall 0.608 0.629 0.500
Table 1: Quadratic Weighted Kappa Scores Across Models and Error Types
The overall IAA scores for the three systems in our study (0.5-0.629) indicate consistent human evaluation. Prior work, including 8d20e0b8-89d8-3d65-bcf5-8c19d56ec4ab, considers kappa scores in the range of 0.41-0.60 to indicate a moderate level of agreement and 0.61-0.80 to indicate substantial agreement. These results confirm the reliability of our annotation process and provide insights for further error analyses studies.
@@ 5.3 Qualitative Error Analysis
5.3.1 Analysis of TRM Errors
@@
“The weather is wonderful today; we should go out for a walk”
@@
“The weather is excellent today; we should go out for a walk!”
@@
“Today’s weather is crazy; we should go out for a walk”
@@ 6 Conclusion
This paper introduces the Ara-HOPE framework as a human-centric approach for evaluating DA-MSA translation, successfully fulfilling its intended objectives through a specialized error taxonomy, an efficient annotation workflow, and a comparative evaluation of different MT systems. The five-category error classification system effectively captures translation challenges unique to DA, while the decision tree protocol improves annotation consistency. Quantitative findings reveal significant differences in performance among Arabic-centric (Jais), general-purpose (GPT-3.5), and baseline (NLLB-200) systems, with dialect-specific terminology and semantic preservation identified as key challenges. By systematically addressing the complexities of DA-MSA translation assessment through rigorous human evaluation, Ara-HOPE establishes reproducible standards for Arabic MT assessment and provides actionable insights to guide future MT systems development.
@@ 7 Limitations
In this work, we used only zero-shot prompting to generate translations. Future research could explore human evaluation of translations produced using alternative prompting strategies, such as few-shot or chain-of-thought prompting. Additionally, the human annotation process was time-consuming. Future work could consider using LLM-as-a-judge approaches to partially or fully automate the annotation process.
Figure 5: Structured decision tree guiding annotators through error classification for evaluating DA-MSA translations using the Ara-HOPE framework.

### v2

@@ 4.3 Questionnaire Design
A structured Excel sheet was created to support annotation and data analysis. Each sheet contained three sets of columns, one per translation engine. Each annotator evaluated 600 translations in total (200 for each translation engine). Figure 1 shows two annotated examples for the Jais system. The first three columns show the DA sentence, the MSA gold translation, and the proposed machine translation. Columns 4-8 represent the five Ara-HOPE error categories, with annotators assigning severity scores from 0 (no error) to 2 (major error). Empty cells indicate a score of 0, and the final column sums the error scores for each sentence.
Figure 1: Example layout of the annotation questionnaire used for evaluating DA-MSA translations across fluency, meaning transfer, and adaptation error categories. To avoid unfairly penalizing systems that preserved meaning but made adaptation errors, adaptation scores were weighted half in the segment total error score calculation.
@@
Three physical copies of the Excel sheet were printed (one for each system) as annotators preferred working on paper. Each contained the same DA-MSA pairs, differing only in the translation engine output. Annotators marked scores manually, and data were later transferred to the digital Excel template.
@@ 5.1 Inter Annotator Agreement
Inter-Annotator Agreement (IAA) measures consistency between evaluators, with high scores indicating a reliable and reproducible annotation framework. In our study, we employed Quadratic Weighted Cohen’s Kappa (QWK) because it effectively handles ordinal severity ratings by assigning greater penalties to larger disagreements, whereas, the ordinary Cohen’s κ treats all disagreements equally, unweighted κ, e.g. 0 vs 1 = 0 vs 2. In addition, quadratic weights are standard for ordinal ratings Cohen (1968). This makes it particularly suitable for error severity judgment on our 0-2 scale, where the magnitude of disagreement matters. The IAA scores are presented in Table 1.
As shown in Figure 6, Meaning Transfer comprises three fine-grained error types (PRN, TRM, and GSMIS). Accordingly, for Meaning Transfer IAA, we compute inter-annotator agreement on an aggregated severity score obtained by summing PRN, TRM, and GSMIS. This reflects annotator agreement on overall meaning-transfer impact rather than on individual sub-error categories. Similarly, SEGS represents the segment-level total error score (produced by summing all five error types’ scores), and the corresponding IAA reflects agreement on overall segment-level error severity.
The IAA scores varied meaningfully across error categories: Fluency showed strong agreement for most MT systems, likely due annotators’ Arabic language expertise, which minimized inaccurate judgments. Meaning Transfer achieved robust agreement, likely due to the clear criteria regarding preservation of original meaning, as established in the Ara-HOPE taxonomy decision tree (Figure 6). Those definitions were particularly refined before the full evaluation during our pilot testing and feedback sessions. This consistency indicates a shared understanding of meaning preservation requirements among our annotators. Adaptation yielded the lowest agreement (0.122-0.28) due to its inherently subjective nature.
A translation has an adaptation error when it seems unnatural or inappropriate to the context, situation, or intent of the source. Unlike fluency or meaning-transfer errors, which can often be localized to specific parts of the translation, adaptation errors require more holistic, and usually subjective judgment. One example is the dialectal sentence:
عم بحكي من واقع تجربة كانت رح تقضي ع حلم حياتي بس الحمد لله الحب اقوى ("I am speaking from experience. It almost destroyed my life’s dream, but thank God, love is stronger"), which the Jais model translated as
أنا أتحدث بناءً على تجربتي الشخصية التي كادت تدمر حلمي الأكبر، ولكن بحمد الله، أثبت الحب أنه أقوى بكثير. (”I speak from my personal experience, which almost destroyed my biggest dream, but thank God, love proved to be much stronger.”)
While both annotators agreed that the translation is fluent and preserves the source meaning, one judged it to be slightly less natural in MSA (ADP = 1, minor adaptation error), whereas the other annotator found it acceptable and contextually appropriate.
The annotator disagreement likely stems from the rendering of phrases suchas بس الحمد لله الحب اقوى (“But thank God, love is stronger”) which was translated as ولكن بحمد الله، أثبت الحب أنه أقوى بكثير. (”But thank God, love proved to be much stronger”), which may be seen as stylistically over-emphatic or slightly less natural in MSA for this context.
@@
SEGS 0.608 0.629 0.500
Table 1: Quadratic Weighted Kappa (QWK) scores across models and error types. Meaning Transfer IAA is computed on aggregated PRN+TRM+GSMIS severity; SEGS reflects total segment-level error severity.
The IAA scores for the three systems in our study on the segment-level total error score (SEGS) was 0.5 to 0.629, indicating reasonably consistent human evaluation.
Prior work, including Landis and Koch (1977), considers kappa scores in the range of 0.41-0.60 to indicate a moderate level of agreement and 0.61-0.80 to indicate substantial agreement. Nevertheless, it is important to note that the interpretation of standard Kappa and Quadratic Weighted Kappa varies considerably.
Aggregating fine-grained error categories into a composite severity score reduces sparsity and stabilizes the marginal distributions, which is known to improve reliability of composite measures Fleiss et al. (2003). Moreover, Quadratic Weighted Kappa assigns smaller penalties to near disagreements Cohen (1968), so cross-category disagreements often translate into small ordinal differences after aggregation. Consequently, agreement computed on aggregated segment-level scores can exceed that of individual categories.
@@ 5.3 Qualitative Error Analysis
5.3.1 Analysis of Dialectal Terminology Errors
@@
“The weather is wonderful today; we should go out for a walk.”
@@
“The weather is excellent today; we should go out for a walk.”
@@
“Today’s weather is crazy; we should go out for a walk.”
@@ 6 Conclusion
This paper introduces the Ara-HOPE framework as a human-centric approach for evaluating DA-MSA translation, successfully fulfilling its intended objectives through a specialized error taxonomy, an efficient annotation workflow, and a comparative evaluation of different MT systems. The five-category error classification system effectively captures translation challenges unique to DA, while the decision tree protocol improves annotation consistency. Quantitative findings reveal significant differences in performance among Arabic-centric (Jais), general-purpose (GPT-3.5), and baseline (NLLB-200) systems, with dialect-specific terminology and semantic preservation identified as key challenges. By systematically addressing the complexities of DA-MSA translation assessment through rigorous human evaluation, Ara-HOPE establishes reproducible standards for Arabic MT assessment and provides actionable insights to guide future MT systems development  [^2] This work is in line with our DA-MSA MT work at Alabdullah et al. (2025) where we examined LLM prompting vs finetuning for Levantine, Egyptian, and Gulf dialects to MSA translation..
@@ 7 Limitations
In this work, we only used zero-shot prompting to generate translations. Future research could explore human evaluation of translations produced using alternative prompting strategies, such as few-shot or chain-of-thought prompting. Additionally, the human annotation process was time-consuming. Future work could consider using LLM-as-a-judge approaches to partially or fully automate the annotation process. Our work focused on human evaluation, and we did not investigate the correlation between human judgment and automatic evaluation metrics, lexicon-based and neural-embedding based, like BLEU and BERTScore. We leave that for future work.
Acknowledgment
We would like to thank our volunteer human annotators, Heba Sado Jazea and Rania Mohammad Daher, for their invaluable time and effort throughout the annotation process. Their careful annotations and insightful feedback were essential to establishing our Ara-HOPE framework. We also thank the VarDial reviewers for the valuable comments.